OpenStudy (anonymous):

slope -4; y intercept (0,-3/5)

OpenStudy (anonymous):

can anyone help me out

OpenStudy (angela210793):

y=-4x-3/5

OpenStudy (anonymous):

y-(-3.5)=-4(x-0) y+3.5=-4x y= -(4x+3.5)